On average, over hundred thousand new businesses are started in the Ghana every year. An article in the
Journal of Business Venturing (Vol. 11, 2017) reported on the activities of entrepreneurs during the
organization creation process. Among the questions investigated were what activities and how many activities
do entrepreneurs initiate in attempting to establish a new business. A total of 71 entrepreneurs were
interviewed and divided into three groups: those that were successful in founding a new firm (34), those still
actively trying to establish a firm (21), and those who tried to start a new firm but eventually gave up (16).
The total number of activities undertaken (i.e., developed a business plan, sought funding, looked for
facilities, etc.) by each group over a specified time period during organization creation was measured and
the ANOVA table has been produced with the sum of squares between groups is 128.70, the Sum of squares
within groups is 27,124.52 with a of 0.013.

(a) Complete the ANOVA table above.
(b). State the null and alternative hypothesis
(c). What is the value of the computed test statistic?
(d) What are the degrees of freedom associated with the test statistic used in this study?
(e) Based on the results of the table above, what can you conclude about the activities of entrepreneurs during
the organization creation process?
Answer #1

(a)

The ANOVA Table is completed as follows:

Source of Variaton Sum of Squares Degrees of Freedom Mean squares F
the sum of squares between groups 128.70 k - 1 = 3 - 1 = 2 128.70/2 = 64.35 64.35/398.89 = 0.1613
the Sum of squares
within groups
27,124.52 70 - 2 = 68 27,124.52/68= 398.89
Total 27,253.22 N - 1 = 71 - 1 = 70

(b)

H0: Null Hypothesis: \mu 1=\mu 2=\mu 3 (There is no significant difference between the means of the activities of entrepreneurs during the organization creation process ) (Claim)

HA: Alternative Hypothesis: \mu 1\neq \mu 2\neq \mu 3 (At least onemean is different from other 2means) (There is a significant difference between the means of the activities of entrepreneurs during the organization creation process )

(c)

the value of the computed test statistic = F = 0.1613

(d)

Degrees of freedom for numerator = 2

Degrees of freedom for denominator = 68

(e)

By Technology,

p value = 0.8514

Since p value = 0.8514 is greater than \alpha = 0.013, the difference is not significant. Fail to reject null hypothesis.

Conclusion:
The data support the claim that there is no significant difference between the means of the activities of entrepreneurs during the organization creation process.